2-(50) Fimd 20,o,는 end the region of operation 2002 3V 2000
Answer #1

Given, ibric {zoon IMA ww + BV ie=ibtic tic VEZ0 200 sie VB=0:740-0.7V loop KVL 7 3=200( ibtic) tiblim) to:7 We have íc=Bib=(& 2:3= ib(1M)+200 (201); > ib= 2.21 MA ic=Bib=200ib alic=0.442 ma ie=ibtic fie= 0.44421 mA Ve=3-200( ibrich 71VC= 2,911 V VCE